OpenStudy (wild_child117):

Solve for a. c = 8 π a bc = 8 π a b 8 π b c = a8 π b c = a c8 π=ac8 π=a c8=ac8=a c8 π b=a

OpenStudy (wild_child117):

wait let me re do this

OpenStudy (the.lit.owl_23):

?

OpenStudy (the.lit.owl_23):

um... ok so you nolonger neet the answer or ...help

OpenStudy (the.lit.owl_23):

ok :)

OpenStudy (ben00):

know that pie = 3.14

OpenStudy (wild_child117):

@brooke..help00

OpenStudy (wild_child117):

yes

OpenStudy (wild_child117):

@perl

OpenStudy (wild_child117):

@mathmale

OpenStudy (mathmale):

Looks like you copied this expression one time too many. Double check the original problem and then share the correct version. c = 8 π a bc = 8 π a b

OpenStudy (wild_child117):

can you view the snip?

OpenStudy (the.lit.owl_23):

omg @Ben00

OpenStudy (mathmale):

Looks as tho' you hit the PASTE button twice: c = 8 π a bc = 8 π a b Fix this, so that we can progress towards finding a in terms of the other variables.

OpenStudy (the.lit.owl_23):

@Wild_Child117

OpenStudy (the.lit.owl_23):

well i gtg soo byeee :) :P

OpenStudy (wild_child117):

oh wait

OpenStudy (wild_child117):

Solve for a. c = 8 π a b A. 8 π b c = a B. c/8π = a C. c/8 = a D. c/8πb = a

OpenStudy (mathmale):

You've typed the same equation twice. This is NOT correct. Solve c=8 pi a b for a. How would you go about isolating 'a' on one side of this equation?

OpenStudy (wild_child117):

Thats how they put it

OpenStudy (wild_child117):

and the equation you typed there isnt a A on the end

OpenStudy (mathmale):

I beg your pardon, but that is NOT "how they put it." Solve for a. c = 8 π a bc = 8 π a b You have typed c = 8 pi a b twice. Focus on Solve for a. c = 8 π a b. How would you isolate 'a' on one side of this equation?

OpenStudy (wild_child117):

idk

OpenStudy (mathmale):

Take c = 8 pi a b and divide both sides by 8 pi b. Type in your result here.

OpenStudy (wild_child117):

8 pi b/ c = 8 pi a b /8 pi b ?

OpenStudy (wild_child117):

it was D

OpenStudy (mathmale):

I asked y ou to take c = 8 pi a b and divide both sides by 8 pi b. Instead, you took 8 pi b and divided that by c. Again: take c = 8 pi a b and divide both sides by 8 pi b, to obtain a formula for 'a'.\[\frac{ c }{ 8 \pi b }=\frac{ 8 \pi a b }{ 8 \pi b }\]

OpenStudy (mathmale):

Simplify this.